OpenStudy (anonymous):

The polynomial x 3 + 5x 2 -57x -189 expresses the volume, in cubic inches, of a shipping box, and the width is (x+3) in. If the width of the box is 15 in., what are the other two dimensions?

OpenStudy (anonymous):

divide x 3 + 5x 2 -57x -189 by (x+3) to get the other factors which represent other dimensions and then use x=15 to find other dimensions

OpenStudy (anonymous):

@matricked can you walk me through please

OpenStudy (anonymous):

@JoannaBlackwelder do you know how to do this?

OpenStudy (triciaal):

as outlined by @matricked except use x = 12 width x + 3 = 15 so x = 15 - 3 = 12

OpenStudy (triciaal):

regular synthetic division then factor the quadratic use x = 12 to find the other dimensions

OpenStudy (anonymous):

@triciaal I'm getting frustrated. I AM SO CONFUSED!!

OpenStudy (triciaal):